# Tony Lindeberg1    Zahra Babaiee2 and Peyman M. Kiasari2

### v1

@@ 2.3.3 Responses to monomials for the 8 “master key filters” operators
Due to spurious variations in the background of the learned filters, these spurious variations may lead to substantial biases in the mean-based and variance-based spread measures, because those spurious variations are weighted with either a linear or a quadratic factor in the first- and second-order moments that define the spatial spread measures. This situation is, thus, in contrast to previous use of such spatial spread measures in Lindeberg (2024, 2025) for characterizing properties of idealized mathematical models of receptive fields with regard to variabilities in their spatial extent under variations of the spatial scale parameter.
@@ 2.5.1 Weighted spatial spread measures
To reduce this substantial source to bias, we will in this section replace the previous uniformly weighted first- and second-order moments by weighted first- and second-order moments, with the weight functions determined from discrete approximations of Gaussian functions with similar values of the scale parameters as we would obtain by estimating the scale parameters of idealized filter models from the spatial variances of the corresponding discrete kernels. Thus, we define weighted spatial mean values and spatial variance of continuous or discrete filter kernels $h(\cdot,\cdot)$ according to:
@@
Specifically, if the continuous receptive field $h(\cdot)$ would be equal to the continuous weighting filter $g(\cdot)$, then the weighted spatial mean value and the weighted spatial variance would correspond to mean value and standard deviation in $L_{2}$-norm. Similarly, if the discrete receptive field $h(\cdot)$ would be equal to the discrete weighting filter $T(\cdot)$, then the weighted spatial mean value and the weighted spatial variance would correspond to mean value and standard deviation in $l_{2}$-norm. For our purpose, we do other hand formulate a hybrid approach, where we use the desired shape of the filter model as a prior to suppress the influence of spurious responses in the learned filter spatial positions in the peripheries of the support regions of the learned filters.
@@
  • For Filter 8 we will either use anisotropic continuous Cartesian-aligned anisotropic Gaussian kernels
@@
    $g_{x}(\cdot,\cdot;\;\sigma_{x},\sigma_{y})=g_{\scriptsize\mbox{1D},x}(x;\;\sigma_{x})\,g_{\scriptsize\mbox{1D}}(y;\;\sigma_{y})$ (83)
@@
    $(\Delta_{x}T)(\cdot,\cdot;\;\sigma_{x},\sigma_{y})=(\Delta_{x}T_{\scriptsize\mbox{1D}})(x;\;\sigma_{x})\,T_{\scriptsize\mbox{1D}}(y;\;\sigma_{y})$ (84)
@@
    $g_{y}(\cdot,\cdot;\;\sigma_{x},\sigma_{y})=g_{\scriptsize\mbox{1D}}(x;\;\sigma_{x})\,g_{\scriptsize\mbox{1D},y}(y;\;\sigma_{y})$ (85)
@@
    $(\Delta_{y}T)(\cdot,\cdot;\;\sigma_{x},\sigma_{y})=T_{\scriptsize\mbox{1D}}(x;\;\sigma_{x})\,(\Delta_{y}T_{\scriptsize\mbox{1D}})(y;\;\sigma_{y})$ (86)
    for a appropriately matched difference operators $\Delta_{y}\in\{\delta_{y+},\delta_{y-},\delta_{y}\}$ in the vertical $y$-direction, with the scale parameters $\sigma_{x}$ and $\sigma_{y}$ determined from variance-based spatial spread measures of the corresponding filters.
The motivation for these choices is to use qualitatively similar shapes for the spatial weighting functions in the spatial spread measures as the absolute values of the corresponding idealized models of receptive field shapes.
@@ 2.6 Estimating variance parameters of receptive fields from weighted spatial spread measures
When using non-uniform weighting kernels when computing the spatial spread measures, the resulting variance-based measures will, however, be affected by the spatial extent of the weighting kernels. In the continuous case, we specifically have the following relations:
@@
Filter $M_{T(\cdot,\cdot)}(|h_{i}(\cdot,\cdot))|$ $V_{T(\cdot,\cdot)}(|h_{i}(\cdot,\cdot))|$
@@
Table 6: The result of computing the discrete weighted spatial mean value $M_{T(\cdot,\cdot)}(|h(\cdot,\cdot))|$ according to Equation (79) and the discrete weighted spatial variance $V_{T(\cdot,\cdot)}(|h(\cdot,\cdot))|$ according to Equation (80) for the normalized versions of the 8 “master key filters” according to (63)–(68) and (73) when using matching shapes of the spatial weighting filters $T(\cdot,\cdot)$ to the qualitative shapes of the “master key filters” according to Section 2.5.2.
@@
As can be seen from a comparison with the corresponding non-weighted spatial spread measures in Tables 1 and 4, the use of spatial weighting in the computation of the variance-based spread measures lead to substantially lower estimates of the corresponding scale parameters, and thus to a substantial decrease of the previous bias, caused by spurious variations in the peripheries of the learned filters.
@@ 3 Modelling the 8 “master key filters” in terms of idealized scale-space operations
In the following, we will use the modelling and analysis results from the previous section, for modelling the 8 “master key filters” in terms of idealized scale-space operations. Specifically, we will explore the implications of using different types of modelling criteria and analyze how they give rise to different sets of scale parameters in the idealized filter models. Overall, we will either consider separate modelling of the different filters of the following forms:
@@
while here (i) restricting the spatial covariance matrix $\Sigma$ in the affine Gaussian kernel $g(x,y;\;s\Sigma)$ to a diagonal matrix according to $\Sigma_{i}=\operatorname{diag}(\sigma_{x,i}^{2},\sigma_{y,i}^{2})$ for each filter $i$ and (ii) restricting the directional derivative operators $\partial_{\varphi}$ and $\partial_{\bot\varphi}$ to Cartesian partial derivatives $\partial_{x}$ and $\partial_{y}$, since the spatial shapes of the learned receptive fields appear to be very much aligned with the Cartesian coordinate directions.
@@
One could, however, also conceive using other discrete approximations of Gaussian derivative operators, such as those in (Lindeberg 2024, 2025)
In the following Sections 3.13.6 we will consider different ways of determining the scale parameters $\sigma_{x,i}$ and $\sigma_{y,i}$ for $i\in\{1,2,3,4,5,6,8\}$, while we will then in Section 3.7 consider a different methodology with one more parameter for determining the parameters $\gamma_{7}$ and $\sigma_{7}$ in an idealized model of Filter 7.
@@
In the visualizations for the resulting Methods A–B to be presented, we will, however, not develop specific variance-based methods for Filter 7, since the idealized model of that filter also comprises an amplitude scaling factor $\gamma_{7}$ for the contribution from the negative Laplacian-of-the-Gaussian. For this reason, we will for Methods A–B instead interweave visualizations of idealized models of Filter 7 based on the either $l_{2}$-norm-based or $l_{1}$-norm-based Methods D2 and C2, to be described in Section 3.7.
Filter $\hat{\sigma}_{x,i}$ $\hat{\sigma}_{y,i}$
1 0.580 0.558
2 0.559 0.580
3 0.617 0.601
4 0.642 0.424
5 1.075 0.800
6 0.771 1.003
8 0.552 0.545
Table 7: Estimates of the horizontal and vertical scale parameters $\hat{\sigma}_{x,i}$ and $\hat{\sigma}_{y,i}$ computed according to Method A, based on estimates from continuous Gaussian derivatives according to (103)–(116) in idealized models of the “master key filters” according to (93)–(100), in turn based on weighted measurements of the variance of the absolute values of the learned filters according to Table 6.
@@
A straightforward way of estimating the scale parameters $\sigma_{x,i}$ and $\sigma_{y,i}$ in the idealized models (93)–(100) is by making use of the closed-form expressions for the weighted variances of the corresponding continuous Gaussian derivative kernels according to (88), (90) and (92). Denoting the weighted variance matrix for each continuous kernel by
@@
we then obtain the following estimates
$\begin{split}\hat{\sigma}_{x,1}=\sqrt{2\,v_{xx,1}},\end{split}$ (103)
$\begin{split}\hat{\sigma}_{y,1}=\sqrt{2\,v_{yy,1}/3},\end{split}$ (104)
$\begin{split}\hat{\sigma}_{x,2}=\sqrt{2\,v_{xx,2}/3},\end{split}$ (105)
$\begin{split}\hat{\sigma}_{y,2}=\sqrt{2\,v_{yy,2}},\end{split}$ (106)
$\begin{split}\hat{\sigma}_{x,3}=\sqrt{2\,v_{xx,3}},\end{split}$ (107)
$\begin{split}\hat{\sigma}_{y,3}=\sqrt{2\,v_{yy,3}/3},\end{split}$ (108)
$\begin{split}\hat{\sigma}_{x,4}=\sqrt{2\,v_{xx,4}/3},\end{split}$ (109)
$\begin{split}\hat{\sigma}_{y,4}=\sqrt{2\,v_{yy,4}},\end{split}$ (110)
$\begin{split}\hat{\sigma}_{x,5}=\sqrt{2\,v_{xx,5}/3},\end{split}$ (111)
$\begin{split}\hat{\sigma}_{y,5}=\sqrt{2\,v_{yy,5}},\end{split}$ (112)
$\begin{split}\hat{\sigma}_{x,6}=\sqrt{2\,v_{xx,6}},\end{split}$ (113)
$\begin{split}\hat{\sigma}_{y,6}=\sqrt{2\,v_{yy,6}/3},\end{split}$ (114)
$\begin{split}\hat{\sigma}_{x,8}=\sqrt{2\,v_{xx,8}},\end{split}$ (115)
$\begin{split}\hat{\sigma}_{y,8}=\sqrt{2\,v_{yy,8}},\end{split}$ (116)
with numerical values of these entities shown in Table 7 based on the numerical values of the variances in Table 6.
Figure 3 shows visualizations of the corresponding idealized receptive fields.
@@
Figure 6: Formal specification of the criteria used for determining the scale estimates $\hat{\sigma}_{x,i}$ and $\hat{\sigma}_{y,i}$ in the horizontal and vertical directions, by matching the diagonal elements of the discrete variance-based spatial spread measures between an idealized receptive field model $h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{x,1},\sigma_{y,1})$ according to (93)–(131) and (100) to the corresponding discrete variance-based spatial spread measures for the normalized versions $h_{i,\scriptsize\mbox{norm}}(\cdot,\cdot)|)$ of the “master key filters” according to (63)–(68) and (73). The parameter $\sigma_{0}$ specifies the spatial extent if the discrete weighting kernel used for computing the weighted variance-based spatial spread measures, and is provided as prior information, chosen as $\sigma_{0}=1$ for all the filters. (Note that since the idealized receptive field models $h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{x,1},\sigma_{y,1})$ are separable with respect to the Cartesian coordinate directions $x$ and $y$, the vertical parameter $\sigma_{y,1}$ does influence the estimate of the horizontal diagonal element $V_{xx,|h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{0},\sigma_{0})|}(|h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{x,1},\sigma_{y,1})|)$. Similarly, the horizontal parameter $\sigma_{x,1}$ does not influence the estimate of the vertical diagonal element $V_{yy,|h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{0},\sigma_{0})|}(|h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{x,1},\sigma_{y,1})|)$.)
Idealized 1(B) Idealized 2(B) Idealized 3(B) Idealized 4(B)
Refer to caption Refer to caption Refer to caption Refer to caption
Idealized 5(B) Idealized 6(B) Idealized 7(C2) Idealized 8(B)
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 7: Idealized models of the 8 “master key filters” in Figure 2, in terms of combinations pure difference operators with spatial smoothing filters based on the discrete analogue of the Gaussian kernel, with the amount of spatial smoothing determined from the spatial extents of the 8 “master key filters” using Method B in Section 3.2, based on requiring the horizontal vs. the vertical discrete weighted variance-based spatial spread measures for the idealized receptive field models to be equal to the weighted discrete spatial spread measures for the corresponding learned filters. (Idealized 1): Non-centered vertical difference operator $(\delta_{y+}T)(m,n;\;\sigma_{x,1},\sigma_{y,1})$. (Idealized 2): Non-centered horizontal difference operator $(\delta_{x-}T)(m,n;\;\sigma_{x,2},\sigma_{y,2})$. (Idealized 3): Non-centered vertical difference operator $(\delta_{y-}T)(m,n;\;\sigma_{x,3},\sigma_{y,3})$. (Idealized 4): Non-centered horizontal difference operator $(\delta_{x+}T)(m,n;\;\sigma_{x,4},\sigma_{y,4})$. (Idealized 5): Centered horizontal difference operator $(\delta_{x}T)(m,n;\;\sigma_{x,5},\sigma_{y,5})$. (Idealized 6): Centered vertical difference operator $(\delta_{y}T)(m,n;\;\sigma_{x,6},\sigma_{y,6})$. (Idealized 7): Sharpening operation $1-\gamma_{7}\,(\nabla_{5}^{2}T)(m,n;\;\sigma_{7},\sigma_{7})$. (Idealized 8): Pure spatial smoothing $T(m,n;\;\sigma_{x,8},\sigma_{y,8})$. (The numerical values of the scale parameters $\sigma_{x,i}$ and $\sigma_{y,i}$ for $i\in\{1,2,3,4,5,6,8\}$ have been determined according to Table 8 and with $\sigma_{7}$ additionally according to Table 13.) (Note that the contrasts of Filters 2 and 3 are reversed in relation to the sign conventions for corresponding “master key filters”.)
@@
A possible limitation of the above computed estimates of the spatial scale parameters from weighted variances of the discrete filters, is that the relationship between the weighted variances of the filters is based on a continuous model in terms of Gaussian derivatives (Equations (88), (90) and (92) for continuous Gaussian derivatives, that lead to the estimates in Equations (103)–(116)), while the filters that we model are genuinely discrete, and specifically corresponding to rather fine scale levels, where discretization effects in relation to the discrete vs. continuous models may therefore be significant.

### v2

@@ 3.3.3 Responses to monomials for the 8 “master key filters” operators
As previously explained towards the end of Section 3.2, that due to the spurious variations in the background of the learned filters, there may be substantial biases in the mean-based and variance-based spread measures, because of these spurious variations being weighted by either a linear or a quadratic factor in the first- and second-order moments that define the spatial spread measures. This situation is, thus, in contrast to previous use of such spatial spread measures in Lindeberg (2024, 2025) for characterizing the properties of idealized mathematical models of receptive fields with regard to variabilities in their spatial extent under variations of the spatial scale parameter.
The underlying reason to this problem is that the idealized filter models, that we are to replace the learned filters with, do all tend to zero further away from the origin, whereas the learned filters contain distinctly non-zero values for image positions where the corresponding idealized filters with similar qualitative shape would assume values very close to zero. Thus, without compensating for this effect, a direct matching of idealized filter shapes to the learned filters based on the regular spatial spread measures in Equations (1)–(4) may overestimate the spatial size of the idealized filters that are to replace the learned filters.
The underlying reason for this is that the spurious variations in the background are all transformed to their absolute values, when computing the spatial spread measures. Specifically, regarding the variance-based spatial spread measure, these absolute values are weighted with the squared distance from the origin. Thereby, such spurious variations may contribute in a different way to the spatial spread measures than their effects in the learned filters, where spurious variations of both positive and negative sign could be expected to in some statistical sense counteract each other.
@@ 3.5.1 Weighted spatial spread measures
To reduce this potentially substantial source of bias, we will in this section replace the previous uniformly weighted first- and second-order moments by weighted first- and second-order moments. For this purpose, we will use weight functions determined from discrete approximations of Gaussian functions with similar values of the scale parameters, as we would obtain by estimating the scale parameters of idealized filter models from the spatial variances of the corresponding discrete kernels. Thus, we define weighted spatial mean values and spatial variance of continuous or discrete filter kernels $h(\cdot,\cdot)$ according to:
@@
Specifically, if the continuous receptive field $h(\cdot)$ would be equal to the continuous weighting filter $g(\cdot)$, then the weighted spatial mean value and the weighted spatial variance would correspond to the mean value and the standard deviation in $L_{2}$-norm. Similarly, if the discrete receptive field $h(\cdot)$ would be equal to the discrete weighting filter $T(\cdot)$, then the weighted spatial mean value and the weighted spatial variance would correspond to the mean value and the standard deviation in $l_{2}$-norm. For our purpose, we do on the other hand formulate a hybrid approach, where we use the desired shape of the filter model as a prior to suppress the influence of spurious responses in the learned filter spatial positions in the peripheries of the support regions of the learned filters.
@@
  • For Filter 8, we will either use anisotropic continuous Cartesian-aligned anisotropic Gaussian kernels
@@
    $g_{x}(x,y;\;\sigma_{x},\sigma_{y})=g_{\scriptsize\mbox{1D},x}(x;\;\sigma_{x})\,g_{\scriptsize\mbox{1D}}(y;\;\sigma_{y})$ (83)
@@
    $(\Delta_{x}T)(m,n;\;\sigma_{x},\sigma_{y})=(\Delta_{x}T_{\scriptsize\mbox{1D}})(m;\;\sigma_{x})\,T_{\scriptsize\mbox{1D}}(n;\;\sigma_{y})$ (84)
@@
    $g_{y}(x,y;\;\sigma_{x},\sigma_{y})=g_{\scriptsize\mbox{1D}}(x;\;\sigma_{x})\,g_{\scriptsize\mbox{1D},y}(y;\;\sigma_{y})$ (85)
@@
    $(\Delta_{y}T)(m,n;\;\sigma_{x},\sigma_{y})=T_{\scriptsize\mbox{1D}}(m;\;\sigma_{x})\,(\Delta_{y}T_{\scriptsize\mbox{1D}})(n;\;\sigma_{y})$ (86)
    for appropriately matched difference operators $\Delta_{y}\in\{\delta_{y+},\delta_{y-},\delta_{y}\}$ in the vertical $y$-direction, with the scale parameters $\sigma_{x}$ and $\sigma_{y}$ determined from variance-based spatial spread measures of the corresponding filters.
  • For Filter 7, we will use either a continuous isotropic [^7] The reason for using an isotropic Laplacian operator here is mainly to obtain a minimum number of parameters in the model. More generally, one could also conceive replacing the isotropic Gaussian kernel in this model with a separable anisotropic Gaussian kernel, that makes use of different scale parameters in the horizontal and the vertical directions. Possibly, one could also consider extending such a model by making use of different scaling constants $\gamma_{x}$ and $\gamma_{y}$, that weigh the contributions from the terms $\partial_{xx}(g(x;\;\sigma_{x}g(y;\;\sigma_{y}))$ and $\partial_{yy}(g(x;\;\sigma_{x}g(y;\;\sigma_{y}))$, respectively, from the horizontal vs. the vertical directions. The discrete sharpening filter (88) could possibly also be extended to corresponding anisotropic image operations. We do, however, leave the possible investigations of such design options regarding the idealized filter models to future work. sharpening filter of the form
    $T_{\text{sharp}}(x,y;\;\sigma,\gamma)=1-\gamma\,(\nabla^{2}g)(x,y;\;\sigma,\sigma),$ (87)
    for an appropriately determined scale parameter $\sigma$ with an appropriately determined scaling constant $\gamma$, where $\nabla^{2}=\partial_{xx}+\partial_{yy}$ denotes the isotropic Laplacian operator, or a discrete isotropic sharpening filter of the form
    $T_{\text{sharp}}(x,y;\;\sigma,\gamma)=\\ =1-\gamma_{7}\,\nabla_{5}^{2}(T_{\scriptsize\mbox{1D}}(m;\;\sigma)\,T_{\scriptsize\mbox{1D}}(n;\;\sigma)),$ (88)
    where $\nabla_{5}^{2}=\delta_{xx}+\delta_{yy}$ denotes the 5-point discrete approximation of the Laplacian operator.
@@ 3.6 Estimating variance parameters of receptive fields from weighted spatial spread measures
When using non-uniform weighting kernels for computing the spatial spread measures, the resulting variance-based measures will, however, be affected by the spatial extent of the weighting kernels. In the continuous case, we specifically have the following relations:
@@
Filter $M_{T(\cdot,\cdot)}(|h_{i}(\cdot,\cdot)|)$ $V_{T(\cdot,\cdot)}(|h_{i}(\cdot,\cdot)|)$
@@
Table 6: The result of computing the discrete weighted spatial mean value $M_{T(\cdot,\cdot)}(|h(\cdot,\cdot)|)$ according to Equation (79) and the discrete weighted spatial variance $V_{T(\cdot,\cdot)}(|h(\cdot,\cdot)|)$ according to Equation (80) for the normalized versions of the 8 “master key filters” according to (63)–(68) and (73) when using matching shapes of the spatial weighting filters $T(\cdot,\cdot)$ to the qualitative shapes of the “master key filters” according to Section 3.5.2.
@@
As can be seen from a comparison with the corresponding non-weighted spatial spread measures in Tables 1 and 4, the use of spatial weighting in the computation of the variance-based spread measures leads to substantially lower estimates of the corresponding scale parameters, and thus to a substantial decrease of the previous bias, caused by spurious variations in the peripheries of the learned filters.
@@ 4 Modelling the 8 “master key filters” in terms of idealized scale-space operations
In the following, we will use the modelling and analysis results from the previous section for modelling the 8 “master key filters” in terms of idealized scale-space operations. Specifically, we will explore the implications of using different types of modelling criteria and analyze how they give rise to different sets of scale parameters in the idealized filter models. Overall, we will either consider separate modelling of the different filters of the following forms:
@@
while here (i) restricting the spatial covariance matrix $\Sigma$ in the affine Gaussian kernel $g(x,y;\;s\Sigma)$ to a diagonal matrix according to $\Sigma_{i}=\operatorname{diag}(\sigma_{x,i}^{2},\sigma_{y,i}^{2})$ for each filter $i$, and (ii) restricting the directional derivative operators $\partial_{\varphi}$ and $\partial_{\bot\varphi}$ to Cartesian partial derivatives $\partial_{x}$ and $\partial_{y}$, since the spatial shapes of the learned receptive fields appear to be very much aligned with the Cartesian coordinate directions.
@@
One could, however, also consider using other discrete approximations of Gaussian derivative operators, such as those in (Lindeberg 2024, 2025)
In the following Sections 4.14.6, we will consider different ways of determining the scale parameters $\sigma_{x,i}$ and $\sigma_{y,i}$ for $i\in\{1,2,3,4,5,6,8\}$, while we will then in Section 4.7 consider a different methodology with one more parameter for determining the parameters $\gamma_{7}$ and $\sigma_{7}$ in an idealized model of Filter 7.
@@
In the visualizations of the results for the resulting Methods A–B to be presented, we will, however, not develop specific variance-based methods for Filter 7, since the idealized model of that filter also comprises an amplitude scaling factor $\gamma_{7}$ for the contribution from the negative Laplacian-of-the-Gaussian. For this reason, we will for Methods A–B instead interweave visualizations of idealized models of Filter 7 based on the either $l_{2}$-norm-based or $l_{1}$-norm-based Methods D2 and C2, to be described in Section 4.7.
Method A Method B Method C1 Method C2 Method D1 Method D2
1 0.580 0.558 0.644 0.583 0.360 0.510 0.458 0.491 0.722 0.644
2 0.559 0.580 0.586 0.644 0.555 0.453 0.448 0.581 0.519 0.558
3 0.617 0.601 0.690 0.674 0.701 0.655 0.671 0.483 0.503 0.495
4 0.642 0.424 0.756 0.460 0.563 0.384 0.420 0.500 0.000 0.380
5 1.075 0.800 1.107 0.945 1.309 0.875 1.387 1.300 1.004 1.193
6 0.771 1.003 0.900 0.889 0.973 1.171 1.090 0.984 1.074 1.038
7 0.654 0.675
8 0.552 0.545 0.609 0.601 0.637 0.587 0.611 0.615 0.608 0.612
Table 7: Summary of the scale estimates obtained for the 8 “master key filters” using the different criteria for fitting the idealized receptive field models to the learned filters: (i) Method A based on direct lookup of anisotropic scale estimates from variance-based spatial spread measures, (ii) Method B based on anisotropic matching of discrete spatial spread measures between the idealized receptive field models and the learned filters, (iii) Method C1 based on anisotropic discrete $l_{1}$-norm minimization, (iv) Method C2 based on isotropic discrete $l_{1}$-norm minimization, (v) Method D1 based on anisotropic discrete $l_{2}$-norm minimization and (vi) Method D2 based on isotropic discrete $l_{1}$-norm minimization. These methods have been explored for all the learned filters except for Filter 7, for which instead the two-parameter modelling scheme in Section 4.7, which also determines a scaling constant for the negative Laplacian-of-the-Gaussian filter in the sharpening operation.
@@
A straightforward way of estimating the scale parameters $\sigma_{x,i}$ and $\sigma_{y,i}$ in the idealized models (95)–(102) is by making use of the closed-form expressions for the weighted variances of the corresponding continuous Gaussian derivative kernels according to (90), (92) and (94). If we denote the weighted variance matrix for each continuous kernel by
@@
then we obtain the following estimates
$\hat{\sigma}_{x,1}$ $=\sqrt{2\,v_{xx,1}},$ $\hat{\sigma}_{y,1}$ $=\sqrt{2\,v_{yy,1}/3},$ (105)
$\hat{\sigma}_{x,2}$ $=\sqrt{2\,v_{xx,2}/3},$ $\hat{\sigma}_{y,2}$ $=\sqrt{2\,v_{yy,2}},$ (106)
$\hat{\sigma}_{x,3}$ $=\sqrt{2\,v_{xx,3}},$ $\hat{\sigma}_{y,3}$ $=\sqrt{2\,v_{yy,3}/3},$ (107)
$\hat{\sigma}_{x,4}$ $=\sqrt{2\,v_{xx,4}/3},$ $\hat{\sigma}_{y,4}$ $=\sqrt{2\,v_{yy,4}},$ (108)
$\hat{\sigma}_{x,5}$ $=\sqrt{2\,v_{xx,5}/3},$ $\hat{\sigma}_{y,5}$ $=\sqrt{2\,v_{yy,5}},$ (109)
$\hat{\sigma}_{x,6}$ $=\sqrt{2\,v_{xx,6}},$ $\hat{\sigma}_{y,6}$ $=\sqrt{2\,v_{yy,6}/3},$ (110)
$\hat{\sigma}_{x,8}$ $=\sqrt{2\,v_{xx,8}},$ $\hat{\sigma}_{y,8}$ $=\sqrt{2\,v_{yy,8}},$ (111)
with numerical values of these entities shown in the first major column in Table 7 based on the numerical values of the variances in Table 6.
The 2nd row in Figure 2 shows visualizations of the corresponding idealized receptive fields.
@@
Figure 6: Formal specification of the criteria used for determining the scale estimates $\hat{\sigma}_{x,i}$ and $\hat{\sigma}_{y,i}$ in the horizontal and vertical directions, by matching the diagonal elements of the discrete variance-based spatial spread measures between an idealized receptive field model $h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{x,1},\sigma_{y,1})$ according to (95)–(100) and (102) to the corresponding discrete variance-based spatial spread measures for the normalized versions $h_{i,\scriptsize\mbox{norm}}(\cdot,\cdot)|)$ of the “master key filters” according to (63)–(68) and (73). The parameter $\sigma_{0}$ specifies the spatial extent of the discrete weighting kernel used for computing the weighted variance-based spatial spread measures, and is provided as prior information, chosen as $\sigma_{0}=1$ for all the filters. (Note that since the idealized receptive field models $h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{x,1},\sigma_{y,1})$ are separable with respect to the Cartesian coordinate directions $x$ and $y$, the vertical parameter $\sigma_{y,1}$ does not influence the estimate of the horizontal diagonal element $V_{xx,|h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{0},\sigma_{0})|}(|h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{x,1},\sigma_{y,1})|)$. Similarly, the horizontal parameter $\sigma_{x,1}$ does not influence the estimate of the vertical diagonal element $V_{yy,|h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{0},\sigma_{0})|}(|h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{x,1},\sigma_{y,1})|)$.)
@@
A possible limitation of the above computed estimates of the spatial scale parameters, from weighted variances of the discrete filters, is that the relationship between the weighted variances of the filters is based on a continuous model in terms of Gaussian derivatives, [^8] According to Equations (90), (92) and (94) for continuous Gaussian derivatives, which lead to the estimates in Equations (105)–(111). while the filters that we model are genuinely discrete. Especially, since the corresponding scale levels are quite fine, discretization effects in relation to the discrete vs. continuous models may therefore be significant.